OpenStudy (anonymous):

Write the equation of a line that passes through the points (2, 0) and (4, -6). y = -3x + 6 y = 3x + 2 y = 3x + 4 y = 3x - 4

OpenStudy (texaschic101):

first we need to find the slope using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (2,0) x1 = 2 and y1 = 0 (4,-6) x2 = 4 and y2 = -6 now we will sub our info into the formula.. slope(m) = (-6 - 0) / (4 - 2) m = -6/2 m = -3 now that we have the slope, we will use the point slope formula: y - y1 = m(x - x1) slope(m) = -3 (2,0) x1 = 2 and y1 = 0 now lets sub y - (-0) = -3(x - 2) y + 0 = -3x + 6 y = 3x + 6 any questions ?

OpenStudy (anonymous):

would it be y=- 3x+6

OpenStudy (texaschic101):

oops....yes...sorry about that

OpenStudy (anonymous):

okay thank you!!!!!

OpenStudy (texaschic101):

anytime :)